two identical spheres each with a mass m and a distance r between their centres , exert a gravitational force of magnitude F on each other. The distance between their centres is halved . what is the magnitude of the force that the spheres exert on each other ?


1
Expert's answer
2021-09-25T13:10:24-0400

The initial force:


"F_i=G\frac{Mm}{r^2},\\\space\\
F_f=G\frac{Mm}{(r/2)^2},\\\space\\
\frac{F_f}{F_i}=4\space\rightarrow\space F_f=4F_i."

The force will increase fourfold.